# Accelerating Diffusion Models in Offline RL via Reward-Aware Consistency Trajectory Distillation

### v1

@@ 7 Conclusion
In this work, we address the challenge of diffusion policy acceleration by introducing reward-aware consistency trajectory distillation (RACTD), which predicts high-reward actions in a single denoising step. RACTD uses a pre-trained diffusion teacher model and a separately trained reward model, leveraging the teacher’s ability to capture multi-modal distributions while prioritizing higher-reward modes to generate high-quality samples from suboptimal training data. Its decoupled training approach avoids the complex concurrent optimization of multiple networks and enables the use of a standalone, noise-free reward model. RACTD outperforms previous state-of-the-art by $8.7\%$ while accelerating its diffusion counterparts up to a factor of $142$.
@@ Appendix A Model Architecture
We follow the model architecture used in Chi et al. [2023] and Prasad et al. [2024] and continue to use the 1D temporal CNN layer for our Unet and FiLM layers to process the conditioning information.
@@ A.2 Model Sizes for Gym-MuJoCo
We follow the setup in Janner et al. [2022], where we use Linear layers and Mish layers Misra [2019] for the reward model. Reward model architecture remains the same across all MuJoCo benchmarks. Model parameters are shown below in Table 8.
@@
Appendix B Training Details
Our models are trained on D4RL dataset [Fu et al., 2020], which was released under Apache-2.0 license.
B.1 Noise Scheduler
We follow the setup in Prasad et al. [2024] and use EDM noise scheduler [Karras et al., 2022] for the teacher model. Particularly, discretization bins are chosen to be 80.
Student model used CTM scheduler [Kim et al., 2023] also with discretization bins of 80.
B.2 Reward Discount Factor
We use $\gamma=0.997$ for hopper, $\gamma=0.99$ for walker2d, and $\gamma=0.9$ for halfcheetah, along with an early termination penalty of -100.
B.3 Weight of Different Losses
@@
hopper-medium-replay 1.0 1.0 1.0
@@
Appendix C More Ablations
C.1 walker-medium
@@ Appendix C More Ablations
Reward-Aware student 98.9 ±$\pm 1.6$ $94.5$$\pm 2.6$
C.2 Comparing with fast sampling algorithms for Maze2d
@@
RACTD (ours) 1 $125.7$$\pm 0.6$ 130.8 ±$\pm 1.8$ $143.8$$\pm 0.0$ 133.4
Appendix D The Trade-off between Mode Selection and Sample Diversity
In this section, we include a discussion about the trade-off between mode selection induced by our reward-aware training and the sample diversity of the student. Naturally, favoring selected modes can led to generation with limited sample diversity as summarized in Table 12. This trade-off between sample diversity and sample optimality observed in RACTD is similar to what has been seen in other generative domains (e.g., language model RLHF [Huang et al., 2024], classifier-guided diffusion, conditional image generation), where preference alignment also often reduces sample diversity. In our case, the reward model acts similarly to a classifier or an alignment reward model, guiding the model toward desirable behaviors and sacrificing some of the sample diversity by design.
@@ Appendix D The Trade-off between Mode Selection and Sample Diversity
Appendix E Limitations and Future work

### v2

@@ 7 Conclusion
In this work, we address the challenge of accelerating diffusion planners at sampling by introducing reward-aware consistency trajectory distillation (RACTD), which predicts high-reward action trajectories in a single denoising step. RACTD uses a pre-trained diffusion teacher planner and a separately trained reward model, leveraging the teacher’s ability to capture multi-modal distributions while prioritizing higher-reward modes to generate high-quality action trajectories from suboptimal training data. Its decoupled training approach avoids the complex concurrent optimization of multiple networks and enables the use of a standalone, noise-free reward model. RACTD outperforms previous state-of-the-art by $9.7\%$ while leveraging consistency trajectory distillation to accelerate its diffusion counterparts up to a factor of $142$.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Model Architecture
We follow the model architecture used in Chi et al. (2023) and Prasad et al. (2024) and continue to use the 1D temporal CNN layer for our Unet and FiLM layers to process the conditioning information.
@@ A.2 Model Sizes for Gym-MuJoCo
Appendix B Reward Model Parameters
We follow the setup in Janner et al. (2022), where we use Linear layers and Mish layers Misra (2019) for the reward model. We stack 4 ConvBlocks for the reward model, each with dimension shown in Table 14. Each ConvBlock consists of a Conv1D layer, groupnorm, and a Mish layer. Reward model architecture remains the same across all MuJoCo benchmarks.
@@ Appendix B Reward Model Parameters
Appendix C Training Details
Our models are trained on D4RL dataset (Fu et al., 2020), which was released under Apache-2.0 license.
C.1 Noise Scheduler
We follow the setup in Prasad et al. (2024) and use EDM noise scheduler (Karras et al., 2022) for the teacher model. Particularly, discretization bins are chosen to be 80.
Student model used CTM scheduler (Kim et al., 2023) also with discretization bins of 80.
C.2 Weight of Different Losses
@@
hopper-medium-replay 1.0 1.0 0.8
@@
Appendix D More results
D.1 Expert and random dataset performance for Gym-MuJoCo
We include the performance on expert and random datasets along with a comparison between dataset quality, ranging from random to expert, with another diffusion-based planning method Diffuser (Janner et al., 2022). Under the trajectory modeling setup, the random dataset provides little valid trajectory and the expert dataset will degrade the problem to behavior cloning. Nevertheless, our method outperforms Diffuser in all settings tested.
hopper random medium-replay medium medium-expert expert NFE
RACTD 31.8±0.0 104.9±2.1 121.0±0.5 129.0±1.3 137.1±0.1 1
Diffuser 6.7±0.1 70.6±1.6 79.6±0.6 106.9±0.2 110.3±0.1 20
Table 10: Performance of Diffuser and RACTD on Gym-MuJoCo random and expert dataset.
D.2 Expert dataset performance for Franka Kitchen
Since kitchen-complete consists of pure expert demonstrations, the problem degrades to behavior cloning and reward guidance becomes less effective.
Table 11: (Offline RL: FrankaKitchen) Performance and sampling efficiency (NFE) of RACTD and diffusion based baselines on FrankaKitchen complete. Each cell has two values: the first is offline model selection and the second (in brackets) is online model selection.
[…]
Diffusion QL 84.0 $\pm 7.4$ ($84.5$ $\pm 6.1$) 55
Consistency AC $51.9$ $\pm 6.0$ ($67.6$ $\pm 2.7$) 22
Consistency BC $45.2$ $\pm 5.0$ ($50.9$ $\pm 3.6$) 22
RACTD(Ours) $\underline{56.3}$ $\pm 8.2$ ($58.1$ $\pm 8.3$) 1
Appendix E More Ablations
E.1 reward objective for walker-medium
@@ Appendix E More Ablations
Reward-Aware student 118.8 $\pm 0.3$ $94.5$ $\pm 2.6$
E.2 Comparing with fast sampling algorithms for Maze2d
@@
RACTD (ours) 1 $125.7$ $\pm 0.6$ 130.8 $\pm 1.8$ $143.8$ $\pm 0.0$ 133.4
E.3 Effect of Reward Objective Weights
The reward loss weight is a hyperparameter in our pipeline that impacts both training stability and performance. We plot the reward curve achieved over 200 epochs of student training with reward loss weights ranging from $[0.3,0.8,1.5]$ on MuJoCo hopper-medium-replay in Figure 4. The mean and standard error are reported across 20 rollouts from intermediate checkpoints.
Figure 4: Ablation on reward objective weight.
E.4 DSM and Reward loss weights
We include the ablation for different DSM and reward loss weights. The other two losses are fixed at optimal when altering the target loss.
DSM weight 0.0 0.3 0.6 0.8 1.0 1.2 1.5 2.0
RACTD $1.5\pm 0.0$ $23.8\pm 1.0$ $35.6\pm 2.2$ $105.2\pm 1.8$ $\textbf{108.3}\pm 1.1$ $100.0\pm 2.3$ $104.9\pm 1.5$ $93.6\pm 3.4$
Table 14: DSM loss weight ablation for RACTD in MuJoCo.
Reward weight 0.0 0.3 0.5 0.7 0.8 0.9 1.0 1.5 2.0
RACTD $50.8\pm 0.3$ $69.8\pm 2.5$ $91.5\pm 2.5$ $\textbf{108.4}\pm 1.4$ $108.3\pm 1.1$ $85.4\pm 3.2$ $100.9\pm 2.6$ $50.5\pm 2.1$ $17.0\pm 0.0$
Table 15: Reward loss weight ablation for RACTD in MuJoCo.
E.5 Number of Sampling Steps
Since our student model is trained for anytime-to-anytime jumps, it naturally extends to multi-step denoising without additional training. Following the approach in Song et al. (2023), given intermediate denoising timesteps $0<t_{1}<t_{2}<T$, we first denoise from $T$ to 0 as usual. We then add noise again to $t_{1}$ and denoise it back to 0, and repeat this process for $t_{2}$. This iterative refinement can enhance generation quality. We evaluate the student using 2, 3, and 4 denoising steps as reported in Table 16. Results show that multi-step sampling barely improves model performance.
NFE Time(s) Score
1 0.0147 $109.5$ $\pm 0.3$
2 $0.0241$ 109.8 $\pm 0.9$
3 $0.0377$ $108.7$ $\pm 0.1$
4 $0.0517$ $107.9$ $\pm 1.6$
E.6 Observation and Planning Horizon
We compare the performance of incorporating different steps of past observations and predicting different numbers of future actions for MuJoCo walker medium-replay here in Table 17. We can see that the performance can be further improved by conditioning on a larger number of observations and a longer planning horizon.
Table 17: We compare the performance of using different numbers of past observations (h) and future actions (c) on MuJoCo walker medium-replay. Results are presented as the mean and standard error across 100 seeds.
walker medium-replay Future actions c=16 Future actions c=8
Past observation h=1 $105.2$ $\pm 1.8$ $104.3$ $\pm 1.4$
Past observation h=4 108.0 $\pm 1.4$ $103.3$ $\pm 2.2$
Appendix F The Trade-off between Mode Selection and Sample Diversity
In this section, we include a discussion about the trade-off between mode selection induced by our reward-aware training and the sample diversity of the student. Naturally, favoring selected modes can led to generation with limited sample diversity as summarized in Table 18. This trade-off between sample diversity and sample optimality observed in RACTD is similar to what has been seen in other generative domains (e.g., language model RLHF (Huang et al., 2024), classifier-guided diffusion, conditional image generation), where preference alignment also often reduces sample diversity. In our case, the reward model acts similarly to a classifier or an alignment reward model, guiding the model toward desirable behaviors and sacrificing some of the sample diversity by design.
@@ Appendix F The Trade-off between Mode Selection and Sample Diversity
Appendix G Theoretical Insights
Here we provide a proof sketch of how our reward-aware consistency trajectory distillation training resembles off-policy deterministic policy gradient (Silver et al., 2014), thus sharing the well-studied theoretical analysis of these methods for offline RL.
According to CTM (Kim et al., 2023) (Section 3), our student model generates a single-step prediction by
$G_{\theta}(\bm{x}_{T},T,0)=g_{\theta}(\bm{x}_{T},T,0)=\bm{x}_{T}+\int_{T}^{0}\frac{\bm{x}_{u}-\mathop{\mathbb{E}}[\bm{x}|\bm{x}_{u}]}{u}du$
From Taylor expansion, we have $G_{\theta}(\bm{x}_{T},T,0)=\mathop{\mathbb{E}}[\bm{x}_{0}|\bm{x}_{T}]+\mathcal{O}(T)$
Let $\mu_{\theta}=\mathop{\mathbb{E}}[\bm{x}_{0}|\bm{x}_{T}]$. Then the gradient of our reward model is
$\nabla_{\theta}R(G_{\theta}(\bm{x}_{T},T,0))=\nabla_{\theta}R(\mu_{\theta})\nabla_{\theta}\mathop{\mathbb{E}}[\bm{x}_{0}|\bm{x}_{T}]=\nabla_{\theta}R(\mu_{\theta})\nabla_{\theta}\mu_{\theta}$
For off-policy deterministic policy gradient[cite], we have
$\nabla_{\theta}J_{\beta}(\mu_{\theta})\approx\mathop{\mathbb{E}}_{\bm{s}\sim\rho^{\beta}}[\nabla_{\theta}\mu_{\theta}(\bm{s})\nabla_{\bm{a}}Q^{\mu}(\bm{s},\bm{a})|_{\bm{a}=\mu_{\theta}(\bm{s})}]$
, where $\rho^{\beta}$ is the behavior policy.
$\nabla_{\theta}J_{\beta}(\mu_{\theta})\approx\mathop{\mathbb{E}}_{s\sim\rho^{\beta}}[\nabla_{\theta}\mu_{\theta}(s)\nabla_{a}Q^{\mu}(s,a)|_{a=\mu_{\theta}(s)}]$
Our method first samples a state $\bm{s}$ from the offline dataset collected by the behavior policy $\rho^{\beta}$ and asks the student model to predict actions $\bm{a}$ given $\bm{s}$. Then, the reward model takes in $\bm{s}$ and $\bm{a}$ and predicts the corresponding Q value. Therefore $R(\mu_{\theta})$ is equivalent to $Q^{\mu}(\bm{s},\bm{a}|_{\bm{s}\sim\rho^{\beta},\bm{a}=\mu_{\theta}(\bm{s})})$. Thus, our reward-aware student model training resembles off-policy deterministic policy gradient.
Appendix H Limitations and Future work